# Text-Free Multi-domain Graph Pre-training: Toward Graph Foundation Models

### v3

@@ 4.3 Downstream adaptation
Mixing prompt. Compared to the unifying prompt, which aims to leverage the unified multi-domain knowledge, the mixing prompt prioritizes knowledge specific to certain source domains. For the target domain, a closely related source domain is likely to provide more relevant prior knowledge. Consequently, it is crucial to align the target domain with each source domain to a varying extent, prioritizing the most relevant one. Thus, we define the mixing prompt, $\vec{p}_{\text{mix}}\in\mathbb{R}^{\tilde{d}}$, as a learnable aggregation of pre-trained domain tokens $\mathcal{T}_{\text{pre}}$, where each domain token $\mathbf{t}_{S_{i}}\in\mathcal{T}_{\text{pre}}$ can be deemed as a pre-trained “prompt” specific to its corresponding source domain $D_{S_{i}}$. Concretely, let
$\textstyle\vec{p}_{\text{mix}}=\sum_{i=1}^{K}\gamma_{i}\vec{t}_{S_{i}},$ (5)
@@
$\vec{H}=\mathtt{GraphEncoder}(G,\mathbf{p}_{\text{uni}}\odot\tilde{\vec{X}};\Theta_{\text{pre}})+\mathtt{GraphEncoder}(G,\mathbf{p}_{\text{mix}}\odot\tilde{\vec{X}};\Theta_{\text{pre}}),$ (6)
@@
In general, for downstream node or graph classification tasks, $\mathcal{L}_{\text{down}}$ follows the universal task template based on subgraph similarity [21], similar to the pre-training loss $\mathcal{L}_{\text{pre}}$. Consider a labeled training set $\Omega_{\text{down}}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ is either a node or a graph, and $y_{i}\in Y$ is the class label of $x_{i}$ among a set of classes $Y$. The loss is then defined as
$\textstyle\mathcal{L}_{\text{down}}(\vec{H};\vec{p}_{\text{uni}},\Gamma)=-\sum_{(x_{i},y_{i})\in\Omega_{\text{down}}}\ln\frac{\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{{h}}_{x_{i}},\overline{\mathbf{{h}}}_{y_{i}})\right)}{\sum_{y\in Y}\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{{h}}_{x_{i}},\overline{\mathbf{{h}}}_{y})\right)},$ (7)
@@ 5.1 Experimental Setup
Datasets. We conduct experiments on five benchmark datasets. (1) Cora [41], (2) Citeseer [42] and (3) Pubmed [42] are citation graphs from different fields (e.g., computing and biomedical sciences). Each dataset consists of a single graph. Each dataset consists of a single graph, with nodes representing scientific publications and edges denoting citations. In line with previous work [8, 9], we treat the edges as undirected. (4) Photo [43] and (5) Computers [44] are both e-commerce networks from Amazon in different categories (e.g., photography and computer related products). Each dataset comprises a single graph, where nodes represent products and edges signify frequent co-purchases between products. (6) Reddit [45] is a social network, where nodes are post and edges represent interactions such as comments and replies between posts. Note that different citation or e-commerce graphs have distinct features. We present additional details of these datasets in Appendix B.
@@
Setup of multi-domain pre-training. To evaluate the performance of multi-domain pre-training, we focus on the more challenging scenario of generalizing to unseen domains in downstream tasks. For Cora, Citesser, Pubmed, Photos and Computers, we designate each dataset individually as the downstream target domain, while employing the remaining four datasets as the four source domains during multi-domain pre-training. When Reddit is used as the target domain, we leverage all five of the previous datasets as source domains. Further details of multi-domain transfer to unseen and seen domains are provided in Appendix D and  E, respectively.
@@
Setup of downstream tasks. We conduct two types of downstream task, namely, node classification, and graph classification. We set the tasks in an $m$-shot classification, i.e., for each class, we randomly sample $m$ instances (nodes or graphs) as supervision. Note that each dataset only comprises a single graph, which cannot be directly applied to graph classification. Therefore, following previous work [46, 47], we generate a set of graphs by constructing ego-networks centered on the target nodes (i.e., those with class labels) in each dataset. Subsequently, we conduct graph classification across these ego-networks, each is labeled identically to its corresponding central node. Given that the $m$-shot tasks are balanced classification, we evaluate performance using accuracy, consistent with prior studies [35, 48, 49]. Note that we pre-train the graph encoder once for each dataset and then apply the same pre-trained model for all types of downstream tasks.
Baselines. We evaluate the performance of MDGPT against state-of-the-art approaches across five primary categories as outlined below. (1) End-to-end graph neural networks: GCN [8] and GAT [9] utilize neighborhood aggregation to iteratively gather messages from adjacent nodes, training in an end-to-end manner without pre-training. (2) Graph pre-training models: DGI/InfoGraph [^1] Original DGI only operates at the node level, while InfoGraph extends it to the graph level. In our experiments, we apply DGI to node classification, and InfoGraph to graph classification. [19, 50] and GraphCL [20]. follow the “pre-train, fine-tune” paradigm. Specifically, they first pre-train the GNN to exploit the intrinsic attributes of the graphs and then fine-tune the pre-trained model on downstream tasks according to task-specific labels. (3) Graph cross-domain model: Hassani [25] pre-trains a attention-based GNN from both contextual and topological views for cross-domain adaptation. (4) Graph prompting models: GPPT [^2] GPPT is specifically designed for downstream node classification task and is not suitable for graph classification. Therefore, in our experiments, we exclusively use GPPT for node classification. [40], GPF[51] and GraphPrompt [21] fall under this category. They employ a self-supervised pre-training tasks and unify downstream tasks into a general template as the pretext task. Subsequently, a single type of prompt is tuned for downstream adaptation. (5) Multi-domain pre-training models: GCOPE [33] is pre-trained on multi-domain datasets through self-supervised tasks, followed by fine-tuning or prompting for downstream adaptation.
We provide further descriptions of these baselines in Appendix C, and their implementation and configuration details, along with those of our approach MDGPT, in Appendix D.
5.2 Few-shot performance evaluation
@@
Table 1: Accuracy of one-shot node classification, where each column indicates a target domain.
@@
Table 2: Accuracy of one-shot graph classification, where each column indicates a target domain.
@@
One-shot performance. We present the results of one-shot node and graph classification tasks on unseen domains in Tables 1 and 2, respectively. (Additional results on seen domains are presented in Appendix E.) We make the follow observations. (1) MDGPT surpasses all baseline methods across all settings, outperforming the best competitor by 3.5–37.9% on node classification and 4.3–13.9% on graph classification. The results demonstrate its effectiveness in multi-domain pre-training and in downstream adaptation to unseen domains. (2) Another text-free multi-domain pre-training method, GCOPE, significantly lags behind MDGPT. Its suboptimal performance can be attributed to two factors. First, it only performs implicit domain alignment via virtual nodes, without explicitly unifying feature spaces across domains. Second, it overlooks source domain-specific knowledge in downstream adaptation. The results also underscore the importance of our domain tokens in aligning diverse domains and enabling a tailored mixture of source domains. (3) GPPT is only comparable to, if not worse than, other baselines since it is not specifically designed for few-shot learning.
@@
Few-shot performance. To examine the performance of MDGPT with more labeled data, we vary the number of shots in both node and graph classification tasks, and report the results in Fig. 3 and Appendix E, respectively. Our findings are as follows. (1) MDGPT significantly surpasses all baselines in low-shot settings with very limited labeled data (e.g., $m\leq 5$), highlighting the ideal use case for our approach. (2) As the number of shots increases, the performance of all methods improves as anticipated. However, MDGPT continues to perform competitively, if not better.
@@
Figure 3: Impact of number of shots on node classification on three target domains.
5.3 Analyses of model insights
To comprehensively analyze the performance of MDGPT, we conduct two ablation studies and a hyperparameter analysis, as follows. We further analyze parameters efficiency in Appendix E.
Figure 4: Data ablation study with a growing number of source domains.
Figure 5: Sensitivity study of $\tilde{d}$, the aligned feature dimension across domains, on three target domains.
Table 3: Model ablation study on the effects of key components.
@@ 5.3 Analyses of model insights
Hyperparameter analysis. We evaluate the impact of the feature dimension, $\tilde{d}$, after performing domain alignment in Eq. (1). We vary $\tilde{d}$ and report the corresponding performance in Fig. 5. We observe that, for both node and graph classification, as $\tilde{d}$ grows from a low value, the performance improves initially since higher dimensions increases model capacity. However, after reaching a peak (around $\tilde{d}=50$), accuracy starts to gradually decline as $\tilde{d}$ grows further, due to the amplified semantic gap in these feature dimensions across various domains. Based on the above observation, $\tilde{d}=50$ is a robust setting in most datasets and tasks.
6 Conclusions
@@
In this paper, toward building a graph foundation model, we explored multi-domain pre-training on graphs, with the objective of learning a comprehensive range of knowledge from diverse domains, and adapting to unseen domains in downstream applications. Our proposed approach MDGPT leverages a series of domain tokens to unify the semantic spaces of multiple source domains, and extract domain-specific knowledge. Moreover, we introduced dual prompts to align the target domain with pre-trained knowledge, including both unified multi-domain knowledge and a tailored mixture of source domain-specific knowledge. Finally, we conducted extensive experiments on six public datasets, demonstrating that MDGPT significantly outperforms various state-of-the-art baselines.
@@
A Broader Impact and Limitations
Broader impact. Our work integrates multi-domain graphs for pre-training, and further adapts to downstream tasks in different domains, presenting a promising path toward developing a graph foundation model. Specifically, a graph foundation model [7] is any model trained on a wide range of graph data, capable of serving as fundamental building blocks for various downstream graph applications. If realized, graph foundation models would offer enhanced expressive power, improved transferrability, and better applicability to complex graph data and a broad range of tasks in various fields. In particular, a foundation model for molecular graphs would fuel important scientific inquiries in chemistry, biomedicine, and material engineering.
Limitations. In our work, we focus on text-free graphs, which involve a more challenging scenario than text-attributed graphs due to the lack of textual descriptions. However, text-attributed graphs are also prevalent, providing auxiliary knowledge and a gateway to connect with large language models. Therefore, a promising future direction is to integrate both text-free and text-attributed graphs across multiple domains for pre-training, to learn a more comprehensive range of knowledge.
B Further Descriptions of Datasets
In this section, we summarize the datasets in Table 4 and further provide a comprehensive description.
@@ B Further Descriptions of Datasets
(1) Cora [^3] https://relational.fit.cvut.cz/dataset/CORA [41] includes 2,708 computing publications, each assigned to one of seven categories. The citation network has 5,429 links. Each publication in the dataset is depicted by a binary word vector, which shows the presence or absence of words from a dictionary containing 1,433 distinct words.
@@
(2) Citeseer [^4] https://nrvis.com/download/data/labeled/citeseer.zip [42] comprises 3,312 computing publications, each classified into one of six categories, which are different from those in Cora. The citation network includes 4,732 links. Each publication is described by a binary word vector that indicates the presence or absence of words from a dictionary containing 3,703 unique words.
@@
(3) PubMed [^5] https://github.com/kimiyoung/planetoid/raw/master/data [42] is composed of 19,717 biomedical publications from diabetes, each classified into one of three categories. The citation network comprises 44,338 links. Each publication in the dataset is depicted by a TF/IDF weighted word vector from a dictionary, indicating the presence of 500 unique words from the dictionary.
@@
(4) Amazon Photo [^6] https://github.com/shchur/gnn-benchmark/blob/master/data/npz/amazon_electronics_photo.npz [43] comprises 7,487 photography related products, categorized into one of eight categories. The co-purchasing network consists of 119,043 edges, which represent frequently bought-together product relationships. Each product is described by a feature vector derived from its metadata and reviews, and is labeled according to its category.
@@
(5) Amazon Computers [^7] https://github.com/shchur/gnn-benchmark/blob/master/data/npz/amazon_electronics_computers.npz [43] comprises 13,752 computer-related products, categorized into ten groups. The co-purchasing network contains 245,861 edges, illustrating frequently bought-together product relationships. Each product is represented by a feature vector generated from its metadata and reviews, and is labeled according to its specific category.
@@
(6) Reddit [^8] https://data.dgl.ai/dataset/reddit.zip [45] consists of 232,965 posts from the social media platform Reddit, categorized into 41 different subreddits. The interaction network includes 11,606,919 edges, representing interactions such as comments and replies between posts. Each post in the dataset is described by a feature vector derived from its textual content and metadata, consisting of 602 unique words.
C Further Descriptions of Baselines

### v4

@@ 4.3. Downstream adaptation
Mixing prompt. Compared to the unifying prompt for the unified multi-domain knowledge, the mixing prompt prioritizes knowledge specific to certain source domains. For the target domain, a closely related source domain is likely to provide more relevant prior knowledge. Consequently, it is crucial to align the target domain with each source domain to a varying extent, prioritizing the most relevant one. Thus, we define the mixing prompt, $\vec{p}_{\text{mix}}\in\mathbb{R}^{\tilde{d}}$, as a learnable aggregation of pre-trained domain tokens $\mathcal{T}_{\text{pre}}$, where each domain token $\mathbf{t}_{S_{i}}\in\mathcal{T}_{\text{pre}}$ can be deemed as a pre-trained “prompt” specific to its corresponding source domain $D_{S_{i}}$. Concretely, let
(5) $\textstyle\vec{p}_{\text{mix}}=\sum_{i=1}^{K}\gamma_{i}\vec{t}_{S_{i}},$
@@
(6) $\scriptsize\vec{H}=\mathtt{GE}(G,\mathbf{p}_{\text{uni}}\odot\tilde{\vec{X}};\Theta_{\text{pre}})+\mathtt{GE}(G,\mathbf{p}_{\text{mix}}\odot\tilde{\vec{X}};\Theta_{\text{pre}}),$
@@
In general, for downstream node or graph classification tasks, $\mathcal{L}_{\text{down}}$ follows the universal task template based on subgraph similarity (Liu et al., 2023b), similar to the pre-training loss $\mathcal{L}_{\text{pre}}$. Consider a labeled training set $\Omega_{\text{down}}=\{(x_{1},y_{1}),(x_{2},y_{2}),\ldots\}$, where $x_{i}$ is either a node or a graph, and $y_{i}\in Y$ is the class label of $x_{i}$ among a set of classes $Y$. The loss is then defined as
(7) $\textstyle\mathcal{L}_{\text{down}}(\vec{H};\vec{p}_{\text{uni}},\Gamma)=-\sum_{(x_{i},y_{i})\in\Omega_{\text{down}}}\ln\frac{\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{{h}}_{x_{i}},\overline{\mathbf{{h}}}_{y_{i}})\right)}{\sum_{y\in Y}\exp\left(\frac{1}{\tau}\text{sim}(\mathbf{{h}}_{x_{i}},\overline{\mathbf{{h}}}_{y})\right)},$
@@ 5.1. Experimental Setup
Datasets. We conduct experiments on five benchmark datasets. (1) Cora (McCallum et al., 2000), (2) Citeseer (Sen et al., 2008) and (3) Pubmed (Sen et al., 2008) are citation graphs from different fields (e.g., computing and biomedical sciences). Each dataset consists of a single graph. Each dataset consists of a single graph, with nodes representing scientific publications and edges denoting citations. In line with previous work (Kipf and Welling, 2017; Veličković et al., 2018a), we treat the edges as undirected. (4) Photo (Shchur et al., 2018) and (5) Computers (McAuley et al., 2015) are both e-commerce networks from Amazon in different categories (e.g., photography and computer related products). Each dataset comprises a single graph, where nodes represent products and edges signify frequent co-purchases between products. (6) Reddit (Hamilton et al., 2017) is a social network, where nodes are post and edges represent interactions such as comments and replies between posts. Note that different citation or e-commerce graphs have distinct features. We present additional details of datasets in Appendix A.
@@
Setup of multi-domain pre-training. To evaluate the performance of multi-domain pre-training, we focus on the more challenging scenario of generalizing to unseen domains in downstream tasks. For Cora, Citesser, Pubmed, Photos and Computers, we designate each dataset individually as the downstream target domain, while employing the remaining four datasets as the four source domains during multi-domain pre-training. When Reddit is used as the target domain, we leverage all five of the previous datasets as source domains. Details of multi-domain transfer to unseen and seen domains are provided in Appendix C and Sect. 5.6, respectively.
@@
Setup of downstream tasks. We conduct two types of downstream task, namely, node classification, and graph classification. We set the tasks in an $m$-shot classification, i.e., for each class, we randomly sample $m$ instances (nodes or graphs) as supervision. Note that each dataset only comprises a single graph, which cannot be directly applied to graph classification. Therefore, following previous work (Lu et al., 2021; Yu et al., 2024b), we generate a set of graphs by constructing ego-networks centered on the target nodes (i.e., those with class labels) in each dataset. Subsequently, we conduct graph classification across these ego-networks, each is labeled identically to its corresponding central node. Given that the $m$-shot tasks are balanced classification, we evaluate performance using accuracy, consistent with prior studies (Yu et al., 2023b; Wang et al., 2020; Liu et al., 2021). Note that we pre-train the graph encoder once for each dataset and then apply the same pre-trained model for all types of downstream tasks.
@@
Table 1. Accuracy of one-shot node classification, where each column indicates a target domain.
@@
Table 2. Accuracy of one-shot graph classification, where each column indicates a target domain.
@@
Baselines. We evaluate the performance of MDGPT against state-of-the-art approaches across five primary categories as outlined below. (1) End-to-end graph neural networks: GCN (Kipf and Welling, 2017) and GAT (Veličković et al., 2018a) utilize neighborhood aggregation to iteratively gather messages from adjacent nodes, training in an end-to-end manner without pre-training. (2) Graph pre-training models: DGI/InfoGraph [^1] Original DGI only operates at the node level, while InfoGraph extends it to the graph level. In our experiments, we apply DGI to node classification, and InfoGraph to graph classification. (Veličković et al., 2018b; Sun et al., 2019) and GraphCL (You et al., 2020). follow the “pre-train, fine-tune” paradigm. Specifically, they first pre-train the GNN to exploit the intrinsic attributes of the graphs and then fine-tune the pre-trained model on downstream tasks according to task-specific labels. (3) Graph cross-domain model: Hassani (Hassani, 2022) pre-trains a attention-based GNN from both contextual and topological views for cross-domain adaptation. (4) Graph prompting models: GPPT [^2] GPPT is specifically designed for downstream node classification task and is not suitable for graph classification. Therefore, in our experiments, we exclusively use GPPT for node classification. (Sun et al., 2022), GPF(Fang et al., 2024) and GraphPrompt (Liu et al., 2023b) fall under this category. They employ a self-supervised pre-training tasks and unify downstream tasks into a general template as the pretext task. Subsequently, a single type of prompt is tuned for downstream adaptation. (5) Multi-domain pre-training models: GCOPE (Zhao et al., 2024) is pre-trained on multi-domain datasets through self-supervised tasks, followed by fine-tuning or prompting for downstream adaptation.
We provide further descriptions of these baselines in Appendix B, and their implementation and configuration details, along with those of our approach MDGPT, in Appendix C.
5.2. Few-shot performance evaluation
@@
One-shot performance. We present the results of one-shot node and graph classification tasks on unseen domains in Tables 1 and 2, respectively. (Additional results on seen domains are presented in Sect. 5.6.) We make the follow observations. (1) MDGPT surpasses all baseline methods across all settings, outperforming the best competitor by 3.5–37.9% on node classification and 4.3–13.9% on graph classification. The results demonstrate its effectiveness in multi-domain pre-training and in downstream adaptation to unseen domains. (2) Another text-free multi-domain pre-training method, GCOPE, significantly lags behind MDGPT. Its suboptimal performance can be attributed to two factors. First, it only performs implicit domain alignment via virtual nodes, without explicitly unifying feature spaces across domains. Second, it overlooks source domain-specific knowledge in downstream adaptation. The results also underscore the importance of our domain tokens in aligning diverse domains and enabling a tailored mixture of source domains. (3) GPPT is only comparable to, if not worse than, other baselines since it is not specifically designed for few-shot learning.
@@
Few-shot performance. To examine the performance of MDGPT with more labeled data, we vary the number of shots in both node and graph classification tasks, and report the results in Fig. 3. Our findings are as follows. (1) MDGPT significantly surpasses all baselines in low-shot settings with very limited labeled data (e.g., $m\leq 5$), highlighting the ideal use case for our approach. (2) As the number of shots increases, the performance of all methods improves as anticipated. However, MDGPT continues to perform competitively, if not better.
@@
Figure 3. Impact of number of shots on node classification on three target domains.
Table 3. Analysis on homophilic and heterophilic data.
Target domain Source domain Accuracy (%)
GCOPE MDGPT
Pubmed Cora 40.92 45.59
Cora + Citeseer 40.58 46.72
Cora + Citeseer + Wisconsin 43.14 48.51
Wisconsin Texas 27.37 27.75
Texas + Cornell 27.55 27.73
Texas + Cornell + Cora 27.14 29.78
Chameleon Wisconsin 23.14 24.81
Wisconsin + Squirrel 24.13 24.95
Wisconsin + Squirrel + Cora 22.61 25.36
Table 4. Model ablation study on the effects of key components.
@@
5.3. Ablation Study
To comprehensively analyze the performance of MDGPT, we conduct two ablation studies as follows.
Figure 4. Data ablation study with a growing number of source domains.
Figure 5. Sensitivity study of $\tilde{d}$, the aligned feature dimension across domains, on three target domains.
5.4. Hyperparameter Analysis
@@
We evaluate the impact of the feature dimension, $\tilde{d}$, after performing domain alignment in Eq. (1). We vary $\tilde{d}$ and report the corresponding performance in Fig. 5. We observe that, for both node and graph classification, as $\tilde{d}$ grows from a low value, the performance improves initially since higher dimensions increases model capacity. However, after reaching a peak (around $\tilde{d}=50$), accuracy starts to gradually decline as $\tilde{d}$ grows further, due to the amplified semantic gap in these feature dimensions across various domains. Based on the above observation, $\tilde{d}=50$ is a robust setting in most datasets and tasks.
5.5. Homophily Sensitivity
To further evaluate the robustness of MDGPT across homophilic and heterophilic graphs, we conduct cross-domain experiments on homophilic datasets (Cora, Citeseer, Pubmed) and heterophilic datasets (Texas (Pei et al., 2020), Wisconsin (Pei et al., 2020), Cornell (Pei et al., 2020), Squirrel (Rozemberczki et al., 2021), Chameleon (Rozemberczki et al., 2021)). The results are shown in Table 3. We observe that regardless of whether the source or target domain graphs are homophilic or heterophilic, MDGPT effectively leverages multi-domain knowledge and bridges the gap between various domains and consistenely outperforms GCOPE. This further demonstrates the effectiveness of MDGPT.
5.6. Node Classification on Seen Domains
We further conduct one-shot node classification tasks on seen target domains. Specifically, we pre-train on the source domains including Cora, Citeseer, Pubmed, Photo and Computers, and evaluate node classification on a seen target domain, Cora or Citeseer, separately. We illustrate the results of several competitive baselines and MDGPT in Table 5, and observe the same trend as those observed in in Sect. 5.2, showing the robustness of MDGPT in dealing with both seen and unseen domains.
Table 5. Accuracy of one-shot node classification on seen target domains.
Method \ Target domain Cora Citeseer
DGI 33.57 $\pm$ 06.83 32.90 $\pm$ 06.10
GraphCL 38.73 $\pm$ 07.43 32.46 $\pm$ 04.75
GraphPrompt 45.98 $\pm$ 09.43 36.45 $\pm$ 08.38
GCOPE 34.95 $\pm$ 07.45 35.18 $\pm$ 07.76
MDGPT 47.44 $\pm$ 10.03 44.96 $\pm$ 10.12
5.7. Parameters Efficiency
We evaluate the parameter efficiency of MDGPT compared to representative baselines. We present the number of tunable parameters during downstream adaptation in Table 6. For GCN, as it follows an end-to-end supervised paradigm, all parameters in the model have to be updated, resulting in the most parameters. For GraphCL and GCOPE, we update only the downstream classifier while freezing the pre-trained model, significantly reducing the number of tunable parameters. For GraphPrompt, since it only tunes a light-wight prompt containing fewer parameters than a classifier, the number of parameters is further reduced. Finally, MDGPT is the most parameter-efficient adaptation, as only the dual prompts are updated. Specifically, the unifying prompt has the same dimension as the unified features, which is generally smaller than the original feature dimension, and the mixing prompt’s dimension is same as the number of source domains, which is a quite small compared to the feature dimension. Despite having the fewest parameters, MDGPT significantly outperforms all baselines, further demonstrating the effectiveness of our design.
Table 6. Comparison of the number of tunable parameters during the downstream adaptation stage.
Method \  Target domain Cora Citeseer Computers Reddit
GCN 14,592 14,336 15,360 23,296
GraphCL 1,792 1,536 2,560 10,496
GraphPrompt 256 256 256 256
GCOPE 1,792 1,536 2,560 10,496
MDGPT 56 56 56 57
6. Conclusions and Future Work
@@
In this paper, we explored multi-domain pre-training on graphs, with the objective of learning a comprehensive range of knowledge from diverse domains, and adapting to unseen domains in downstream applications. Our proposed approach MDGPT leverages a series of domain tokens to unify the semantic spaces of multiple source domains, and extract domain-specific knowledge. Moreover, we introduced dual prompts to align the target domain with pre-trained knowledge, including both unified multi-domain knowledge and a tailored mixture of source domain-specific knowledge. Finally, we conducted extensive experiments on six public datasets, demonstrating that MDGPT significantly outperforms various state-of-the-art baselines.
In this work, we focused on text-free graphs, which are more challenging than text-attributed graphs due to the lack of textual descriptions. However, text-attributed graphs are also important, providing auxiliary knowledge and a gateway to connect with large language models. Therefore, a promising future direction is to integrate both text-free and text-attributed graphs across multiple domains for pre-training, in order to learn a more comprehensive range of knowledge.
@@
A. Further Descriptions of Datasets
In this section, we provide a comprehensive description of datasets.
@@ A. Further Descriptions of Datasets
(1) Cora [^3] https://relational.fit.cvut.cz/dataset/CORA (McCallum et al., 2000) includes 2,708 computing publications, each assigned to one of seven categories. The citation network has 5,429 links. Each publication in the dataset is depicted by a binary word vector, which shows the presence or absence of words from a dictionary containing 1,433 distinct words.
@@
(2) Citeseer [^4] https://nrvis.com/download/data/labeled/citeseer.zip (Sen et al., 2008) comprises 3,312 computing publications, each classified into one of six categories, which are different from those in Cora. The citation network includes 4,732 links. Each publication is described by a binary word vector that indicates the presence or absence of words from a dictionary containing 3,703 unique words.
@@
(3) PubMed [^5] https://github.com/kimiyoung/planetoid/raw/master/data (Sen et al., 2008) is composed of 19,717 biomedical publications from diabetes, each classified into one of three categories. The citation network comprises 44,338 links. Each publication in the dataset is depicted by a TF/IDF weighted word vector from a dictionary, indicating the presence of 500 unique words from the dictionary.
@@
(4) Amazon Photo [^6] https://github.com/shchur/gnn-benchmark/blob/master/data/npz/amazon_electronics_photo.npz (Shchur et al., 2018) comprises 7,487 photography related products, categorized into one of eight categories. The co-purchasing network consists of 119,043 edges, which represent frequently bought-together product relationships. Each product is described by a feature vector derived from its metadata and reviews, and is labeled according to its category.
@@
(5) Amazon Computers [^7] https://github.com/shchur/gnn-benchmark/blob/master/data/npz/amazon_electronics_computers.npz (Shchur et al., 2018) comprises 13,752 computer-related products, categorized into ten groups. The co-purchasing network contains 245,861 edges, illustrating frequently bought-together product relationships. Each product is represented by a feature vector generated from its metadata and reviews, and is labeled according to its specific category.
@@
(6) Reddit [^8] https://data.dgl.ai/dataset/reddit.zip (Hamilton et al., 2017) consists of 232,965 posts from the social media platform Reddit, categorized into 41 different subreddits. The interaction network includes 11,606,919 edges, representing interactions such as comments and replies between posts. Each post in the dataset is described by a feature vector derived from its textual content and metadata, consisting of 602 unique words.
B. Further Descriptions of Baselines